# WEATHER-5K: A Large-scale Global Station Weather Dataset Towards Comprehensive Time-series Forecasting Benchmark

### v1

@@ 4.1 Problem Definition
Considering $N$ weather stations around the world and each station collects $V$ meteorological variables, the data of all weather stations can be represented by a spatial-temporal time-series $X\in\mathbb{R}^{N\times T\times V}$ for a given look-back window of fixed length $T$. At timestamp $t$, time-series forecasting is to predict $\hat{X}_{t+1:t+\tau}=\{X_{t+1},...,X_{t+\tau}\}$ based on the past $T$ frames $X_{t-T+1:t}=\{X_{t-T+1},...,X_{t}\}$. Here, $\tau$ is the length of the forecast horizon. Using $X$ and $\hat{X}$ to represent the observation data and the forecasted data, respectively, the process of GSWF can be simplified by a mapping: $\hat{X}=\mathcal{M}(X)$, where $\mathcal{M}$ can be different kinds of time-series forecasting methods. For example, by setting $N=1$ and ignoring the spatial information, many state-of-the-art time-series forecasting methods [49, 34, 35, 13, 15, 41, 51, 37, 16] can be explored on this task. When $N$ is multiple scattered stations, method [12] based on spatial-temporal modeling can also be applied to this task.
@@ 4.2 Evaluation Metrics
Overall performance. Mean Absolute Error (MAE) and Mean Square Error (MSE) are used to evaluate the overall performance of the GSWF. MAE measures the predictive robustness of an algorithm but is insensitive to outliers, whereas MSE is sensitive to outliers and can amplify errors.
Extreme performance. In the field of weather forecasting, an important evaluation is the ability to predict extreme weather events, such as extreme high or low temperatures. However, MAE and MSE alone do not adequately capture this ability. Therefore, we introduce a specialized metric, Symmetric Extremal Dependency Index (SEDI) [52, 53] to evaluate the performance of extreme weather forecasting for time-series forecasting models. SEDI can assess the model’s capability to predict extreme weather conditions. It classifies each prediction in its station location as either extreme or normal weather based on upper quantile thresholds ($90\%,95\%,98\%$, and $99.5\%$) or lower quantile thresholds ($10\%,5\%,2\%$, and $0.5\%$). The calculation of SEDI value can be formulated as:
@@
where $\hat{X}<Q^{p}_{lower}$ and $X<Q^{p}_{lower}$ judge whether the predicted or observed data point belongs to the extreme event or not based on the threshold $Q^{p}_{lower}$, and vice versa for the upper percentiles. The reason behind setting two opposite percentiles is that extreme small value and extreme large value are both crucial for weather forecasting. For example, the smaller value in temperature may represent the winter storm while the larger value may represent the heatwave. $\mathrm{SEDI}\in[0,1]$ quantifies the model’s ability to correctly identify extreme weather events. A higher SEDI value indicates better performance in extreme weather prediction.
@@ 4.3 Experimental Protocols
Dataset splitting. The WEATHER-5K dataset is divided into three subsets: training, validation, and testing datasets. The training set consists of weather data from 2014 to 2021, the validation set includes data from the year 2022, and the testing set comprises data from the year 2023. The division of the dataset follows an 8:1:1 ratio, allowing the model to be trained on sufficient historical data, validated on a separate year, and tested on the most recent data for accurate evaluation.
@@
Table 2: Forecasting results on our WEATHER-5K. We report the results at 4 different prediction lengths: 24, 72, 120, and 168, where the input length is 48. The selective Baselines are developed for specific research, such as “Electricity (E)", “Exchange (X)", “Influenza (I)", “Traffic (T)", “Weather (W)" and “General (G)" domains.
@@
Baselines Lead Time Temperature Dewpoint Wind Rate Wind Direc. Sea Level Overall
MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE
Informer 2021 [34] EW 24 1.88 7.51 1.94 8.30 1.30 3.62 60.7 6906.9 2.01 10.56 13.6 1387.4
72 2.75 14.84 2.86 17.24 1.53 4.86 71.5 8251.4 4.24 39.24 16.4 1631.4
120 3.11 18.21 3.25 21.50 1.60 5.38 75.7 8504.5 5.15 54.31 18.3 1720.4
168 3.24 20.24 3.43 24.89 1.63 5.65 76.2 8718.4 5.26 58.42 18.1 1764.4
Autoformer 2021 [13] EXITW 24 1.93 8.64 2.06 9.57 1.42 3.97 66.5 7710.0 2.26 12.78 15.2 1553.4
72 2.72 15.14 2.97 18.38 1.54 5.14 75.4 9111.5 4.25 42.34 17.8 1846.7
120 3.21 20.27 3.34 23.12 1.58 5.73 79.2 9143.5 4.83 48.88 18.1 1868.3
168 3.43 21.71 3.56 22.55 1.64 5.95 79.8 9435.8 5.32 61.85 18.5 1885.7
Pyraformer 2021 [35] EW 24 1.75 6.92 1.83 7.88 1.30 3.58 61.8 6930.2 1.90 9.72 13.7 1391.7
72 2.47 13.03 2.67 15.39 1.52 4.97 72.0 8222.4 3.76 33.67 16.5 1657.9
120 2.77 16.04 3.00 18.95 1.59 5.37 75.1 8610.7 4.43 43.91 17.4 1739.0
168 2.95 17.95 3.20 21.06 1.61 5.56 76.4 8773.5 4.77 49.97 17.8 1773.6
FEDformer 2022 [15] EXITW 24 1.98 8.45 2.02 9.25 1.36 3.91 66.0 7384.1 2.13 11.43 14.7 1483.4
72 2.87 16.50 3.01 18.70 1.59 5.31 76.2 8824.8 4.15 37.60 17.6 1780.6
120 3.19 20.29 3.36 23.10 1.66 5.71 79.0 9143.3 4.81 48.86 18.4 1848.3
168 3.35 22.12 3.54 25.21 1.68 5.88 79.7 9189.2 5.01 53.39 18.7 1859.2
DLinear 2023 [40] EXITW 24 2.71 13.82 2.47 12.36 1.44 4.34 66.6 8234.5 3.09 21.34 15.3 1657.3
72 3.55 23.05 3.48 22.85 1.62 5.37 75.0 9250.8 4.64 45.83 17.7 1869.6
120 3.90 27.60 3.89 27.72 1.67 5.70 77.3 9510.6 5.19 56.22 18.4 1925.6
168 4.11 30.38 4.11 30.58 1.69 5.88 78.4 9630.0 5.48 61.73 18.8 1951.7
PatchTST 2023 [37] EXITW 24 2.05 9.26 2.16 10.58 1.40 4.20 66.2 7765.8 2.19 12.54 14.8 1560.5
72 2.82 16.60 3.06 19.96 1.60 5.39 75.2 9067.8 4.28 42.46 17.4 1830.5
120 3.15 20.32 3.43 24.39 1.66 5.79 77.8 9452.6 5.09 57.29 18.2 1912.1
168 3.33 22.54 3.63 26.94 1.69 6.00 79.0 9638.1 5.51 65.3 18.6 1951.7
Corrformer 2023 [12] W 24 1.99 8.21 2.09 9.47 1.38 3.83 66.7 7832.3 2.19 12.39 14.9 1584.4
72 2.74 15.16 2.99 18.40 1.56 4.91 75.6 9111.7 4.27 42.36 17.8 1846.7
120 3.06 18.63 3.34 22.48 1.61 5.56 78.0 9477.4 5.08 57.13 18.1 1915.8
168 3.09 18.69 3.36 22.53 1.63 5.69 78.9 9636.0 5.34 61.83 18.4 1938.8
Mamba 2023 [41] G 24 1.98 8.59 2.01 9.52 1.37 4.02 66.0 7709.5 2.21 12.73 14.7 1548.9
72 2.79 16.00 2.90 18.11 1.55 5.11 75.1 8863.9 4.29 41.88 17.3 1789.0
120 3.03 18.47 3.18 21.02 1.58 5.28 76.7 8931.2 4.93 52.56 17.9 1805.7
168 3.16 19.88 3.32 22.53 1.59 5.35 77.4 8958.8 5.21 57.37 18.1 1812.8
iTransformer 2024 [16] ETW 24 1.82 7.49 1.93 8.80 1.32 3.77 63.2 7358.8 1.99 10.84 14.1 1478.0
72 2.60 14.46 2.84 17.5 1.52 4.96 73.2 8713.3 4.14 40.65 16.9 1758.2
120 2.97 18.36 3.24 22.16 1.59 5.42 76.4 9192.2 4.95 54.67 17.8 1858.6
168 3.18 20.64 3.48 24.89 1.64 5.67 78.0 9441.1 5.36 62.31 18.3 1910.9
@@
Informer [34] 11.8 49.5 9.2 39.2 2.1 6.7 0.12 2.9 9.8 35.7
Autoformer [13] 12.4 52.1 8.3 38.9 0.3 7.8 0.13 1.6 10.4 32.1
Pyraformer [35] 10.7 54.8 7.2 40.1 0.6 7.2 0.06 1.1 10.5 26.2
FEDformer [15] 11.9 50.9 9.9 40.7 2.9 9.5 0.08 0.7 7.5 21.4
DLinear [40] 5.8 18.8 3.2 19.9 0.3 5.1 0.13 1.7 2.8 17.5
PatchTST[37] 10.9 50.8 8.9 42.4 0.5 8.9 0.10 2.2 13.5 36.7
Corrformer [12] 10.9 48.9 8.4 39.9 1.7 8.4 0.12 0.9 8.9 30.9
Mamba [41] 10.0 51.3 7.5 40.6 0.9 8.1 0.05 1.0 10.1 31.3
iTransformer [16] 14.1 55.0 10.4 44.8 1.3 10.3 0.14 2.3 15.9 37.5
@@
Baselines. As discussed in Section 2.1, many time-series forecasting methods can be modified to fit this task. We here compare 9 baselines. 1) Temporal-only methods: The transformer-based models won the most popularity in the past three years, some methods between 2021 and 2024, like Informer (2021) [34], Autoformer (2021) [13], Pyraformer (2021) [35], FEDformer (2022) [15], PatchTST (2023) [37], iTransformer (2024) [16] are selected to represent the state-of-the-art long-term forecasting methods. Apart from them, we also adopt the efficient MLP-based DLinear (2023)[40] and the new trending architecture Mamba (2023) [41] as baselines. 2) Spatial-temporal method: Temporal-only deep models that do not consider the spatial correlations, we, therefore, survey the Corrformer (2023) [12] as a baseline model that considers the dynamic characteristics of correlations among weather stations at different locations.
@@
Task settings. To facilitate fair comparison between different baselines, we align the input length of all baselines. The final setting is to predict the $\tau$-step future based on 48 historical steps. Specifically, we predict the future weather conditions for lead times of 1, 3, 5, and 7 days, corresponding to predicting the 24-step, 72-step, 120-step, and 168-step future data, respectively. Note the report results are only performed once instead of multiple times in original implementations. This does not affect the comparison as we observe the results are stable for different seeds due to the large dataset volume.
@@
Implementation details. We develop and implement the baselines based on the Time-Series-Library. [^2] https://github.com/thuml/Time-Series-Library To ensure a fair comparison, we made no changes to the intermediate network structures of all baseline models. Instead, we make corresponding adjustments to the inputs and outputs to adapt them to the WEATHER-5K dataset. Additionally, all optimization parameters are kept the same. For instance, the training is performed for a total of 300,000 iterations, starting with a learning rate of 1e-4. We employ the cosine decay strategy and gradually decay the learning rate to 0 by the end of training. The batch size for all models is set to 1,024 except for Correformer. During the validation phase, an early stopping is executed if training loss does not decrease for three consecutive times. The checkpoint with the lowest validation loss encountered prior to the early stop is saved and used for testing. Experiments are conducted on an 224 Intel(R) Xeon(R) Platinum 8480CL CPUs @ 3.80 GHz, 2.0 TB RAM computing server, equipped with 8 NVIDIA H800 GPUs.
4.4 Quantitative Comparison
@@
Table 2 reports the MAE and MSE for five variables under different predictive lengths. Some findings are concluded as follows. The simple linear implementation DLinear [40] shows a relatively poor performance while the early method, like Pyraformer[35], shows a significant advantage among all baselines. Overall, the remaining baselines show similar performance. There could be several reasons for this. Firstly, the existing time-series forecasting models have relatively small parameter settings and computational capacity, which can achieve good fitting performance on small-scale datasets but may not be suitable for large-scale datasets. Additionally, these models, except Corrformer [12], only consider temporal dependencies and overlook the spatial distribution differences and correlations in meteorological data. Furthermore, we observe that Corrformer, despite considering spatial relationships, does not perform as well as some simpler models like Informer [34]. One possible reason for this may be that Corrformer’s spatial modeling relies on pre-defined local regions. Moreover, we noticed that different models exhibit preferences for some variables. For example, Informer [34] performs better in predicting wind speed and direction than other methods. In addition, some models show preferences for long-term forecasting, such as the Mamba [41] model, which achieves the best performance in long-term predictions even though it has a general short-term performance. This suggests that the Mamba structure may be more suitable for long-term forecasting. Finally, based on the current results, the cumulative error stabilizes and increases significantly after the third day. This indicates the predictive errors are large after three days, and there is still ample room for improvement in long-term prediction performance.
@@
Table 3 presents a report on the predictive performance of various models for extreme values. Our finding is that the current time-series forecasting models studied in this paper can not effectively capture the extreme values, especially with the lower and upper quantiles at $0.05\%$ and $99.5\%$. Additionally, it is observed that the performance of wind prediction is the worst among all variables. This can be attributed to the non-stationary nature of the wind distribution, which makes it extremely challenging to predict accurately. The evaluation results also indicate a research direction for future time-series forecasting, which is to pay more attention to extreme values.
4.5 Ablation Study
Figure 3: The influence of the input length.
The ablation experiments are conducted to explore the impact of different input lengths on predictions. Here, we utilized the iTransformer [16] model for verification. Specifically, we set the input length to $24,48,72,96$, and $120$, while keeping the predicted length constant at 72. The experimental results are shown in Figure 3, illustrating the performance variation of MAE for four weather variables as the input sequence length increases. The results demonstrate that the performance for some variables (temperature and wind) is slightly improved when increasing the input length. On the other hand, the MAE error for sea level pressure rises slightly. We ultimately set the input length as 48 to balance computation and performance.
5 Conclusion, Discussion, and Limitation
To facilitate accurate, efficient, and scalable weather forecasting for global weather stations, we introduce WEATHER-5K as a new benchmark dataset. WEATHER-5K encompasses numerous global stations, providing comprehensive, long-term meteorological data. This dataset enables state-of-the-art time-series forecasting methods to be easily adopted and yield promising results. However, we also noticed that current methods might still lag behind numerical weather prediction models, particularly for longer lead times. This means WEATHER-5K would present new challenges and opportunities, fostering advanced techniques and innovative research. We highlight key future research opportunities inspired by this dataset.
  • Developing the spatial-aware time-series forecasting methods. Our observation shows that different regions exhibit distinct weather patterns, such as variations between hemispheres or coastal and inland climates, while current forecasting methods do not fully capture global spatial relationships. By incorporating spatial information, such as weather stations’ geographical coordinates, researchers can leverage the relationships between nearby stations to enhance prediction accuracy. Spatially aware models can more effectively capture local weather patterns and dependencies, resulting in more precise forecasts.
@@ 5 Conclusion, Discussion, and Limitation
  • Briding GSWF with numerical prediction models. We are aware that most Numerical weather prediction (NWP) models [9, 8, 11] can provide robust global atmospheric forecasts. By utilizing outputs from these models, we can develop bias correction models tailored to meteorological stations. Leveraging this diverse information can significantly enhance weather prediction accuracy at these stations.
Currently, our work has some limitations. Firstly, the large size of the dataset increases development costs and computational demands, posing challenges for researchers in terms of resources. Secondly, there is a scarcity of existing research utilizing spatial modeling in time-series forecasting. Consequently, the baselines we implemented do not fully leverage the potential of the WEATHER-5K dataset. We welcome and look forward to future researchers fully utilizing WEATHER-5K to develop more advanced time series forecasting methods.
Appendix A Dataset Documentation

### v2

@@ 4.1 Problem Definition
Considering $N$ weather stations around the world and each station collects $V$ meteorological variables, the data of all weather stations can be represented by a spatial-temporal time-series $X\in\mathbb{R}^{N\times T\times V}$ for a given look-back window of fixed length $T$. At timestamp $t$, time-series forecasting is to predict $\hat{X}_{t+1:t+\tau}=\{X_{t+1},...,X_{t+\tau}\}$ based on the past $T$ frames $X_{t-T+1:t}=\{X_{t-T+1},...,X_{t}\}$. Here, $\tau$ is the length of the forecast horizon. Using $X$ and $\hat{X}$ to represent the observation data and the forecasted data, respectively, the process of GSWF can be simplified by a mapping: $\hat{X}=\mathcal{M}(X)$, where $\mathcal{M}$ can be different kinds of time-series forecasting methods. For example, by setting $N=1$ and ignoring the spatial information, many state-of-the-art time-series forecasting methods Li et al. (2022); Zhou et al. (2021); Liu et al. (2021); Wu et al. (2021); Zhou et al. (2022); Gu & Dao (2023); Wang et al. (2024a); Nie et al. (2022); Liu et al. (2024b) can be explored on this task. When $N$ is multiple scattered stations, method Wu et al. (2023) based on spatial-temporal modeling can also be applied to this task.
@@ 4.2 Evaluation Metrics
Overall performance.
Mean Absolute Error (MAE) and Mean Square Error (MSE) are used to evaluate the overall performance of the GSWF. MAE measures the predictive robustness of an algorithm but is insensitive to outliers, whereas MSE is sensitive to outliers and can amplify errors.
Extreme performance metric.
In addition to standard metrics MAE and MSE, the precision in predicting extreme weather events, such as unusually high or low temperatures, is crucial in real-world applications. Hence, we introduce a specialized metric, the Symmetric Extremal Dependence Index (SEDI). Han et al. (2024b); Xu et al. (2024). It classifies each prediction in its station location as either extreme or normal weather based on upper quantile thresholds ($90\%,95\%,98\%$, and $99.5\%$) or lower quantile thresholds ($10\%,5\%,2\%$, and $0.5\%$). The calculation of SEDI value can be formulated as:
@@
where $\hat{X}<Q^{p}_{lower}$ and $X<Q^{p}_{lower}$ judge whether the predicted or observed data point belongs to the extreme event or not based on the threshold $Q^{p}_{lower}$, and vice versa for the upper percentiles. Here we access both two opposite percentiles for accessing extremely small values (e.g., winter storm) and extremely large values (e.g., heatwave). $\mathrm{SEDI}\in[0,1]$ quantifies the model’s ability to correctly identify extreme weather events. Higher SEDI indicates better performance in extreme weather prediction.
@@ 4.3 Experimental Protocols
The WEATHER-5K dataset is divided into three subsets: training (with years 2014-2021), validation (with yea 2022), and test (with year 2023), which follows an 8:1:1 ratio. allowing the model to be trained on sufficient historical data, validated on a separate year, and tested on the most recent data for accurate evaluation.
@@
Table 2: Benchmarks on our WEATHER-5K. The results are based on 4 different prediction lengths: 24, 72, 120, and 168, where the input length is 48. Those baselines are developed for specific research, such as “Electricity (E)", “Exchange (X)", “Influenza (I)", “Traffic (T)", “Weather (W)" and “General (G)" domains. ECMWF-HRES is the physical-based NWP model, representing the most accurate weather forecasting model. Underline is the second-best performance.
@@
Baselines Lead Time Temperature Dewpoint Wind Rate Wind Direc. Sea Level
MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE
ECMWF-HRES Best NWP Model 24 1.76 7.39 1.85 7.94 1.48 4.53 63.8 7158.3 0.86 2.68
72 1.87 8.01 1.94 8.48 1.52 4.76 72.4 8215.6 1.06 3.31
120 1.99 8.79 2.14 10.87 1.58 5.11 75.4 8647.7 1.38 5.15
168 2.15 10.06 2.32 12.56 1.66 5.59 78.3 8945.7 1.87 9.52
Pyraformer 2021 EW Best TSF Model 24 1.75 6.92 1.83 7.88 1.30 3.58 61.8 6930.2 1.90 9.72
72 2.47 13.03 2.67 15.39 1.52 4.97 72.0 8222.4 3.76 33.67
120 2.77 16.04 3.00 18.95 1.59 5.37 75.1 8610.7 4.43 43.91
168 2.95 17.95 3.20 21.06 1.61 5.56 76.4 8773.5 4.77 49.97
Informer 2021 EW 24 1.88 7.51 1.94 8.30 1.30 3.62 60.7 6906.9 2.01 10.56
72 2.75 14.84 2.86 17.24 1.53 4.86 71.5 8251.4 4.24 39.24
120 3.11 18.21 3.25 21.50 1.60 5.38 75.7 8504.5 5.15 54.31
168 3.24 20.24 3.43 24.89 1.63 5.65 76.2 8718.4 5.26 58.42
Autoformer 2021 EXITW 24 1.93 8.64 2.06 9.57 1.42 3.97 66.5 7710.0 2.26 12.78
72 2.72 15.14 2.97 18.38 1.54 5.14 75.4 9111.5 4.25 42.34
120 3.21 20.27 3.34 23.12 1.58 5.73 79.2 9143.5 4.83 48.88
168 3.43 21.71 3.56 22.55 1.64 5.95 79.8 9435.8 5.32 61.85
FEDformer 2022 EXITW 24 1.98 8.45 2.02 9.25 1.36 3.91 66.0 7384.1 2.13 11.43
72 2.87 16.50 3.01 18.70 1.59 5.31 76.2 8824.8 4.15 37.60
120 3.19 20.29 3.36 23.10 1.66 5.71 79.0 9143.3 4.81 48.86
168 3.35 22.12 3.54 25.21 1.68 5.88 79.7 9189.2 5.01 53.39
DLinear 2023 EXITW 24 2.71 13.82 2.47 12.36 1.44 4.34 66.6 8234.5 3.09 21.34
72 3.55 23.05 3.48 22.85 1.62 5.37 75.0 9250.8 4.64 45.83
120 3.90 27.60 3.89 27.72 1.67 5.70 77.3 9510.6 5.19 56.22
168 4.11 30.38 4.11 30.58 1.69 5.88 78.4 9630.0 5.48 61.73
PatchTST 2023 EXITW 24 2.05 9.26 2.16 10.58 1.40 4.20 66.2 7765.8 2.19 12.54
72 2.82 16.60 3.06 19.96 1.60 5.39 75.2 9067.8 4.28 42.46
120 3.15 20.32 3.43 24.39 1.66 5.79 77.8 9452.6 5.09 57.29
168 3.33 22.54 3.63 26.94 1.69 6.00 79.0 9638.1 5.51 65.3
Corrformer 2023 W 24 1.99 8.21 2.09 9.47 1.38 3.83 66.7 7832.3 2.19 12.39
72 2.74 15.16 2.99 18.40 1.56 4.91 75.6 9111.7 4.27 42.36
120 3.06 18.63 3.34 22.48 1.61 5.56 78.0 9477.4 5.08 57.13
168 3.09 18.69 3.36 22.53 1.63 5.69 78.9 9636.0 5.34 61.83
Mamba 2023 G 24 1.98 8.59 2.01 9.52 1.37 4.02 66.0 7709.5 2.21 12.73
72 2.79 16.00 2.90 18.11 1.55 5.11 75.1 8863.9 4.29 41.88
120 3.03 18.47 3.18 21.02 1.58 5.28 76.7 8931.2 4.93 52.56
168 3.16 19.88 3.32 22.53 1.59 5.35 77.4 8958.8 5.21 57.37
iTransformer 2024 ETW 24 1.82 7.49 1.93 8.80 1.32 3.77 63.2 7358.8 1.99 10.84
72 2.60 14.46 2.84 17.5 1.52 4.96 73.2 8713.3 4.14 40.65
120 2.97 18.36 3.24 22.16 1.59 5.42 76.4 9192.2 4.95 54.67
168 3.18 20.64 3.48 24.89 1.64 5.67 78.0 9441.1 5.36 62.31
@@
ECMWF-HRES 37.4 82.6 35.4 76.4 10.2 40.8 13.1 45.4 77.5 89.7
Informer 11.8 49.5 9.2 39.2 2.1 6.7 0.12 2.9 9.8 35.7
Autoformer 12.4 52.1 8.3 38.9 0.3 7.8 0.13 1.6 10.4 32.1
Pyraformer 10.7 54.8 7.2 40.1 0.6 7.2 0.06 1.1 10.5 26.2
FEDformer 11.9 50.9 9.9 40.7 2.9 9.5 0.08 0.7 7.5 21.4
DLinear 5.8 18.8 3.2 19.9 0.3 5.1 0.13 1.7 2.8 17.5
PatchTST 10.9 50.8 8.9 42.4 0.5 8.9 0.10 2.2 13.5 36.7
Corrformer 10.9 48.9 8.4 39.9 1.7 8.4 0.12 0.9 8.9 30.9
Mamba 10.0 51.3 7.5 40.6 0.9 8.1 0.05 1.0 10.1 31.3
iTransformer 14.1 55.0 10.4 44.8 1.3 10.3 0.14 2.3 15.9 37.5
Baselines.
@@
We here compare 9 baselines as mentioned in Section 2, which can be categorized as: 1) Temporal-only methods, which includes popular transformer-based long-term forecasting models between 2021 and 2024: Informer (2021) Zhou et al. (2021), Autoformer (2021) Wu et al. (2021), Pyraformer (2021) Liu et al. (2021), FEDformer (2022) Zhou et al. (2022), PatchTST (2023) Nie et al. (2022), iTransformer (2024) Liu et al. (2024b), MLP-based models: DLinear (2023)Zeng et al. (2023) and the new trending architecture Mamba (2023) Gu & Dao (2023) as baselines. 2) Spatial-temporal method: We also include a recent method, Corrformer (2023) Wu et al. (2023) as a baseline model that considers the dynamic characteristics of correlations among weather stations at different locations.
@@
Task settings. To facilitate fair comparison between different baselines, we align the input length of all baselines. The final setting is to predict the $\tau$-step future based on 48 historical steps, where the input length is chosen based on the experimental results shown in Figure 4 b), which illustrates the performance variation for four weather variables as the input sequence length increases. To trade off the training cost and performance, we ultimately set the input length as 48 to balance computation and performance. Specifically, we predict the future weather conditions for lead times of 1, 3, 5, and 7 days, corresponding to predicting the 24-step, 72-step, 120-step, and 168-step future data, respectively. Note the report results are only performed once instead of multiple times in original implementations. This does not affect the comparison as we observe the results are stable for different seeds due to the large dataset volume.
@@
Implementation details. We develop and implement the baselines based on the Time-Series-Library (Tim, ). The training is performed for a total of 300,000 iterations, starting with a learning rate of 1e-4. We employ the cosine decay strategy and gradually decay the learning rate to 0 by the end of training. The batch size for all models is set to 1,024 except for Correformer. During the validation phase, an early stopping is executed if training loss does not decrease for three consecutive times. The checkpoint with the lowest validation loss encountered prior to the early stop is saved and used for testing. Experiments are conducted on 224 Intel(R) Xeon(R) Platinum 8480CL CPUs @ 3.80 GHz, 2.0 TB RAM computing server, equipped with 8 NVIDIA H800 GPUs.
4.4 Observations and Findings
RQ1: How do TSF models compare with NWP models in terms of general performance?
@@ 4.4 Observations and Findings
. In long-term predictions, where lead time $\geq 72$ hours, the NWP model ECMWF-HRES (EC, ) is overall superior to all existing TSF methods across almost all variables apart from wind speed and direction, while in the short-term prediction, where lead time $=24$ hours, some TSF methods (Pyraformer Liu et al. (2021)) shows a comparable performance, which shows there is a long way for TSF models to be the applicable model for weather forecasting. Details can be found in Table 2, which reports the MAE and MSE for five variables under different predictive lengths. We also found that the simple linear implementation DLinear Zeng et al. (2023) shows a poor performance while the early method, like Pyraformer Liu et al. (2021), shows a significant advantage among all baselines. Overall, the remaining baselines show similar performance. There could be several reasons for this. Firstly, the existing TSF models have relatively small parameter settings and computational capacity, which can achieve good fitting performance on small-scale datasets but may not be suitable for large-scale datasets. Additionally, these models, except Corrformer Wu et al. (2023), only consider temporal dependencies and overlook the spatial distribution differences and correlations in meteorological data. Furthermore, we observe that Corrformer, despite considering spatial relationships, does not perform as well as some simpler models like Informer Zhou et al. (2021). One possible reason for this may be that Corrformer’s spatial modeling relies on pre-defined local regions. Moreover, we noticed that different models exhibit preferences for some variables. For example, Informer Zhou et al. (2021) performs better in predicting wind speed and direction than other methods. In addition, some models show preferences for long-term forecasting, such as the Mamba Gu & Dao (2023) model, which achieves the best performance in long-term predictions even though it has a general short-term performance. This suggests that the Mamba structure may be more suitable for long-term forecasting. Finally, based on the current results, the cumulative error stabilizes and increases significantly after the third day. This indicates the predictive errors are large after three days, and there is still ample room for improvement in long-term prediction performance.
RQ2: Can TSF models perform well in real-world extreme weather prediction?
@@
Another finding is that current TSF models struggle to predict extreme values, whereas numerical models excel at forecasting these extreme values, which is crucial for extreme weather event assessment. Table 3 presents a report on the predictive performance of various models for extreme values. Our finding is that the current time-series forecasting models studied in this paper can not effectively capture the extreme values, especially with the lower and upper quantiles at $0.05\%$ and $99.5\%$. Additionally, it is observed that the performance of wind prediction is the worst among all variables. This can be attributed to the non-stationary nature of the wind distribution, which makes it extremely challenging to predict accurately. The evaluation results also indicate a research direction for future time-series forecasting, which is to pay more attention to extreme values.
Figure 4: a) Model performance vs complexity. b) The performance impact of input length.
RQ3: Are large-parameter TSF models necessary for weather forecasting?
Highly parameterized models with intensive computational requirements do not necessarily enhance the predictive performance of time series forecasting. Figure 4 compares the model complexity and prediction accuracy (average results over 72 hours) of different baselines. This figure give some guidances on the future direction of TSF models. For instance, blindly scaling model parameters may not necessarily improve prediction accuracy. Pyraformer (Liu et al., 2021), for example, not only outperforms all baselines in terms of performance but also has relatively low training costs and parameter counts. On the other hand, Corrformer (Wu et al., 2023), with its high parameter count, significantly increases training time without bringing about substantial performance improvements. Since time series models are typically applied in specific domains and the data is updated rapidly, a model with fewer parameters is more conducive to practical deployment and iterative updates.
RQ4: How to improve future weather forecasting methods leveraging the advantages of TSF and NWP models?
@@
Bridging the TSF models with NWP models would greatly enhance the performance of GSWF. We are aware that most NWP models Bi et al. (2023); Lam et al. (2023); Han et al. (2024a) can provide robust global atmospheric forecasts. By utilizing outputs from these models, we can develop bias correction models tailored to meteorological stations. Leveraging this diverse information by bridging GSWF with numerical prediction models could potentially enhance weather prediction accuracy at these stations.
5 Conclusion and Limitation
To facilitate accurate, efficient, and scalable weather forecasting for global weather stations, we introduce WEATHER-5K as a new benchmark dataset. WEATHER-5K encompasses numerous global stations, providing comprehensive, long-term meteorological data. This dataset enables state-of-the-art time-series forecasting methods to be easily adopted and yield promising results. However, we also noticed that current methods might still lag behind numerical weather prediction models, particularly for longer lead times. This means WEATHER-5K would present new challenges and opportunities, fostering advanced techniques and innovative research.
Limitation. Currently, we have three major limitations. 1). Spatial Quality: The current version of WEATHER-5K offers observations from a global network, but coverage remains sparse in regions like Africa and South America. Future enhancements will focus on integrating data from diverse sources, such as MADIS reports (including METER and Mesonet), to achieve denser station distribution and improve evaluation accuracy. 2). Missing Data Handling. To maintain data integrity, we utilized interpolated results from the ERA5 dataset to fill in missing observations. While this interpolation introduces some errors, minimally affecting temperature, dew point temperature, and pressure but significantly impacting wind data, we ensured that only a small fraction of missing values were filled in to minimize these errors. 3). Lack of Spatial-Temporal Methods: There is a scarcity of research utilizing spatial modeling in time-series forecasting. Consequently, the baselines we implemented do not fully leverage the potential of the WEATHER-5 dataset. We look forward to future researchers fully utilizing WEATHER-5K to develop more advanced time-series forecasting methods.
Appendix A Appendix
Appendix B Dataset Documentation